OpenStudy (anonymous):

I'll give medals... Write the equation of a quadratic function that has zeros of -5 and 8 in standard form. HELP PLEASE

OpenStudy (anonymous):

(x+5)(x-8) = x^2-3x-40.

OpenStudy (anonymous):

Thought so .. how do you know the "a" value though?

OpenStudy (anonymous):

Do you know how to foil?

OpenStudy (anonymous):

No .. Explain ?

OpenStudy (anonymous):

FOIL = Firsts, Outsides, Insides, Lasts. So. I know that -5 and 8 HAVE to be zeroes of this equation. Quadratics that look like (x...) (x,,,) = 0 means that either the first bunch of numbers (x...) or the second bunch (x,,,)has to be 0, as anything multiplied by 0 = 0. So, (x+5) and (x-8) are zeroes, because (-5+5) = 0 or (8-8) = 0. Now you foil this out to get a quadratic in standard form. Take the first x in (x+5,) and distribute it across the (x-8), giving you x^2 - 8x. Then, take the 5 from (x+5) and distribute it across the (x-8) to get 5x - 40. Combine x^2-8x+5x-40 and combine like terms.

OpenStudy (anonymous):

So, the x*x = Firsts, x*-8 = outsides, 5*x = insides, 5*-8 = lasts.

OpenStudy (anonymous):

Oh okay thanks